OpenStudy (anonymous):

A student says that the functions f(x) = 2x+2 and g(x) =2x-2 are inverse functions because their graphs are parallel. Is the student's reasoning correct? Justify your answer

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

Please help me @Nnesha

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Someone please help me....

jimthompson5910 (jim_thompson5910):

what do you get if you solve y = 2x+2 for x?

OpenStudy (anonymous):

Umh, you would get y=x-2/2

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

And then the other one would be y=x+2/2

jimthompson5910 (jim_thompson5910):

yes \[\Large y = \frac{x-2}{2}\] that is equivalent to \[\Large y = \frac{1}{2}x - 1\] what is the slope of the second equation I wrote?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

or would it be -1? Because I know that would be the y intercept right

jimthompson5910 (jim_thompson5910):

it's 1/2

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

so in y = 2x+2, the slope is 2 in y = (1/2)x - 1, the slope is 1/2

jimthompson5910 (jim_thompson5910):

the slopes are not equal, so the lines aren't parallel

OpenStudy (anonymous):

And that would mean that the two functions are not inverses, right?

jimthompson5910 (jim_thompson5910):

no when we solved for x and swapped x and y, we found the inverse

OpenStudy (anonymous):

But don't inverse functions have to equal x?

jimthompson5910 (jim_thompson5910):

I'm just saying that not all inverses are parallel

OpenStudy (anonymous):

No i know that, but are they in this scenario?

OpenStudy (anonymous):

Are the two equations inverses? I wouldn't think so because they aren't equal to x... Or am I wrong

jimthompson5910 (jim_thompson5910):

oh you mean y=2x+2 and y=2x-2?

OpenStudy (anonymous):

Yeah

jimthompson5910 (jim_thompson5910):

well you just found the inverse of y = 2x+2 was y = (1/2)x-1 so it is NOT y=2x-2

jimthompson5910 (jim_thompson5910):

y=2x+2 and y=2x-2 aren't inverses of each other

OpenStudy (anonymous):

Oh ok. Yeah, that's what I thought, but I just wanted to make sure. Do you think you could help me with another problem?

OpenStudy (anonymous):

Another problem that also has to do with inverse functions...?

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (anonymous):

Find the inverse of each function. Then use the definition of inverse functions to verify that the two functions are inverses. 1. f(x) = -3x+3

OpenStudy (anonymous):

I know that the inverse for this one is f -1(x)=3-x/3

OpenStudy (anonymous):

The other function is g(x)=0.25x+.6

jimthompson5910 (jim_thompson5910):

use parenthesis and say (3-x)/3

jimthompson5910 (jim_thompson5910):

keep in mind that 3-x/3 without parenthesis means \(\LARGE 3 - \frac{x}{3}\)

OpenStudy (anonymous):

Ok. And the inverse for the second one is f-1(x)=-4(.6-x) right?

OpenStudy (anonymous):

Would I have to plug them into each other to see if they are inverses following the definition of inverse functions?

jimthompson5910 (jim_thompson5910):

y=0.25x+0.6 has the inverse y = 4(x-0.6) or y = 4x-2.4

jimthompson5910 (jim_thompson5910):

yeah you need to confirm that f(g(x)) = x and g(f(x)) = x

OpenStudy (anonymous):

ok, so what would I do next?

OpenStudy (anonymous):

Ok. and they are not inverses right?

jimthompson5910 (jim_thompson5910):

which 2?

OpenStudy (anonymous):

there's more than two?

jimthompson5910 (jim_thompson5910):

I'm lost about what you're asking

OpenStudy (anonymous):

You asked me which two

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

I apologize. I just don't understand this

jimthompson5910 (jim_thompson5910):

so you're given 2 functions, right? which 2 functions are you given again?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

Would it be the two inverses of the functions?

jimthompson5910 (jim_thompson5910):

f(x) = -3x+3 and g(x)=0.25x+.6 right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

so if you can show that f(g(x)) = x and g(f(x)) = x are both true, then you have proven they are inverses of each other

OpenStudy (anonymous):

ok. Thank you for your help.. I'll post another question if I need help.